E2T9 Consider a solid sphere with radius R with a charge uniformly spread throughout its volume. The magnitude of the spheres electric field is E1 at a point P1 on the spheres surface and is lE2| at a point P2 at a distance jR from the spheres center. In the relationshiplE,l-blE,L, what is the value of b? A. 0 B. 1 C. 2 D. 4 F. Other
Answer #1


on the surface

magnitude of electric field E1 = k*Q/R^2


at a distance r < R

consider a sphere of radius r < R


volume charge density sigma = Q/((4/3)*pi*R^3)


volume of sphere(r) V = (4/3)*pi*r^3


charge enclosed inside the sphere of radius r = Qin = sigma*V

Qin = Q*(r/R)^3

from gauss law


flux = Qin/e0

E2*A = Qin/eo

E2*4*pi*r^2 = Q*(r/R)^3

E2 = Q*r/(4*pi*e0*R^3) = k*Q*r/R^3


for r = R/2

E2 = k*Q/(2*R^2)


E2 = E1/2

E1 = 2*E2


E = b*E2

therefore b =2 <<< ----------ANSWER


OPTION (C)